Calculate q, w, and Delta U for 10.0 g of ice melting at 0.0 ºC and at a constant pressure of 1.000 atm.

Use the data reported below, but note that other necessary information to solve the problem is not provided and that you are expected to look it up.

Density of Ice = 0.917 g/mL

Specific heat of fusion = 79.7 cal/g

Specific heat of vaporization = 539.6 cal/g

NOTE: I believe that one of the equations that is supposed to be used is q = m * Deltafus H, meaning that the heat of fusion is one of the items that is supposed to be looked up. Also, a lot of sources are giving wrong answers for this question. In order to check our work, we were given what the correct answers are supposed to be. q is either going to equal 3335 or 203.3, and w is either going to equal 0.092, or -15.5. These are not supposed to be provided, so whats needed is the work to find these answers.
Answer #1

i) q = m * \DeltafusH = 10 g * 79.7 cal/g = 797 cal (or) 3335 J (Note: 1 cal = 4.184 J, i.e. 797*4.184 ~ 3335)

ii) Work (w) = -P.V = -1 atm * 10 g/(0.917 g/10-3 L) = -0.0109 L.atm (or) -1.105 J (Note: 1 L.atm = 101.325 J)

iii) \DeltaU = \DeltafusH + w = 6002.4 J/mol - 1.105 J/0.556 mol = 6000.4 J/mol (or) 6.0004 kJ/mol